1. Questions 8 to 12—convert the following number of moles to number of particles:
  2. 2.0 moles of CaCl2
  3. 1.5 moles of copper(II) ions
  4. 0.01 moles of H2O
  5. 5.0 moles of gold atoms
  6. 0.0025 moles of HCl
  7. What is the mass of 1 mole of Mg atoms?
  8. What is the molar mass of aluminum?
  9. What is the mass of 1 mole of oxygen atoms?
  10. What is the mass of 1 mole of oxygen molecules, O2?
  11. Calculate the molar mass of NaOH.
  12. What is the mass of 1.0 mole of CaCl2?
  13. Calculate the molar mass of Cu(NO3)2.
  14. Calculate the molar mass of Al2(SO4)3.
1
Expert's answer
2021-06-09T07:14:08-0400

Solution:

One mole of any substance contains 6.022×1023 particles.

Thus:

2)

Number of CaCl2 molecules = (2.0 mol CaCl2) × (6.022×1023 molecules / 1 mol) = 1.20×1024 molecules

Number of CaCl2 molecules = 1.20×1024 CaCl2 molecules

3)

Number of Cu2+ ions = (1.5 mol Cu2+) × (6.022×1023 ions / 1 mol) = 9.03×1023 Cu2+ ions

Number of Cu2+ ions = 9.03×1023 Cu2+ ions

4)

Number of H2O molecules = (0.01 mol H2O) × (6.022×1023 molecules / 1 mol) = 6.02×1021 molecules

Number of H2O molecules = 6.02×1021 H2O molecules

5)

Number of Au atoms = (5.0 mol Au) × (6.022×1023 atoms / 1 mol) = 3.01×1024 Au atoms

Number of Au atoms = 3.01×1024 Au atoms

6)

Number of HCl molecules = (0.0025 mol HCl) × (6.022×1023 molecules / 1 mol) = 1.51×1021 molecules

Number of HCl molecules = 1.51×1021 HCl molecules

7)

The molar mass of Mg is 24.305 g/mol.

So, 1 mol of Mg atoms has a mass: (1 mol Mg) × (24.305 g Mg / 1 mol Mg) = 24.305 g Mg

8)

The molar mass of Al is 26.98 g/mol.

9)

The molar mass of O is 15.999 g/mol.

So, 1 mol of oxygen atoms has a mass: (1 mol O) × (15.999 g O / 1 mol O) = 15.999 g O atoms

10)

The molar mass of O2 is 31.998 g/mol.

So, 1 mol of O2 molecules has a mass: (1 mol O2) × (31.998 g O2 / 1 mol O2) = 31.998 g O2 molecules

11)

M(NaOH) = Ar(Na) + Ar(O) + Ar(H) = 22.989 + 15.999 + 1.008 = 39.996 (g/mol).

The molar mass of NaOH is 39.996 g/mol.

12)

M(CaCl2) = Ar(Ca) + 2×Ar(Cl) = 40.078 + 2×35.453 = 110.984 (g/mol).

The molar mass of CaCl2 is 110.984 g/mol.

So, 1 mol of CaCl2 has a mass: (1 mol CaCl2) × (110.984 g CaCl2 / 1 mol CaCl2) = 110.984 g CaCl2

13)

M(Cu(NO3)2) = Ar(Cu) + 2×Ar(N) + 6×Ar(O) = 63.546 + 2×14.007 + 6×15.999 = 187.554 (g/mol).

The molar mass of Cu(NO3)2 is 187.554 g/mol.

14)

M(Al2(SO4)3) = 2×Ar(Al) + 3×Ar(S) + 12×Ar(O) = 2×26.98 + 3×32.065 + 12×15.999 = 342.143 (g/mol).

The molar mass of Al2(SO4)3 is 342.143 g/mol.


Answers:

2) 1.20×1024 CaCl2 molecules;

3) 9.03×1023 Cu2+ ions;

4) 6.02×1021 H2O molecules;

5) 3.01×1024 Au atoms;

6) 1.51×1021 HCl molecules;

7) 24.305 g Mg;

8) the molar mass of Al is 26.98 g/mol;

9) 15.999 g O atoms;

10) 31.998 g O2 molecules;

11) the molar mass of NaOH is 39.996 g/mol;

12) 110.984 g CaCl2;

13) the molar mass of Cu(NO3)2 is 187.554 g/mol;

14) the molar mass of Al2(SO4)3 is 342.143 g/mol.